OpenStudy (hihi):

A batch of cookies calls for 3/4 cup of flour. You want to make 1/2 of a batch. How much flour do you need?

OpenStudy (bamboo888):

\[3/4 \div 2 = KCC\] KCC(R)= Keep Change Change Reciprocal Keep the 3/4, change the division sign to multiplication, Change or flip the last number, 2 = 2/1 or two over one and flip it to 1/2 and just multiply across so its: 3/4 x 1/2 = 3/8